# A mix of long-duration hydrogen and thermal storage enables large-scale electrified heating in a renewable European energy system

### v1

@@ 1 Introduction
The direct or indirect electrification of sectors such as transport, industry, and heating is crucial for the complete decarbonization of economies around the world [international_energy_agency_net_2024]. Yet, most studies at the nexus of LDES and weather variability focus only on the power sector and neglect the impact of weather variations on the demand side of a sector-coupled renewable system [gotske_designing_2024]. Accounting for 16% percent of total greenhouse gas emissions in the European Union in 2024 [rozsai_m_jrc-idees-2021_2024], the heating sector takes center stage in the continent’s decarbonization efforts [commission_renewable_2023]. Previous studies have shown that serving a substantial part of residential and commercial heat demand with heat pumps can have challenging consequences for the power system [bloess_power--heat_2018, zeyen_mitigating_2021, roth_power_2023, roth_power_2024]. Furthermore, there is a substantial degree of interannual variation in temperatures, and correspondingly in heat demand and temperature-dependent heat pump efficiencies. There is evidence that cold spells can coincide with times of renewable supply shortages [van_der_wiel_influence_2019, mockert_meteorological_2023] exacerbating system stress.
@@ 2 Results
In the following, we focus primarily on the link between electrification of residential and commercial heat and LDES needs. The Supplemental Information Section SI.4 provides several additional results. We summarize all considered scenarios in Table 1 and provide additional details in Section 4.3.
@@ 2.1 Electrified heat demand leads to significantly higher long-duration electricity storage requirements
In Figure 1, the weather years are ordered by LDES requirement in the Decent scenario. Notably, the No Heat scenario does not reflect the same ordering. Weather years that have high LDES requirements in the scenario without electrified heat do not necessarily rank high when considering electrified heat (cf. 1989/90) and vice versa (cf. 1941/42). Additionally, the distribution of optimal LDES capacities widens substantially. In the No Heat scenario, the standard deviation of total LDES energy capacities across all weather years is 10 TWh. In the scenario Decent, the standard deviation increases to 66 TWh.
@@
Optimal long-duration electricity storage capacities by weather year. Grey dots represent optimal capacities in the No Heat scenario. Dark-red dots represent optimal capacities in the Decent scenario. Light-red crosses represent optimal capacities in the Decent - Mean scenario, in which weather-year specific heat demand is replaced by long-run mean heat demand.
@@ 2.2 Heat demand drives storage needs through ‘leverage’ and ‘compound’ effects
In a third set of scenario runs (Decent - Mean), we replace the time series of heat demand and heat pump coefficients of performance (COP) in a given weather year by their long-run hourly averages across all weather years in the sample. These curves (see Figure SI.11) isolate the inherent seasonality of electricity demand for heating and abstract from cold spells or unusually warm periods. The results of these model runs (light red crosses in Figure 1) highlight the drivers of increasing LDES requirements. Heat demand is seasonal and electrified heating naturally adds most load in winter months. As extreme system states of VRE scarcity mostly occur during winter months [kittel_quantifying_2024, kittel_coping_2025, gotske_designing_2024], the additional load amplifies the energy deficit, triggering additional LDES requirements by functioning as a lever. Across the sample of weather years, this leverage effect accounts for 98 TWh or 75% of the total LDES increase on average.
For some years, the Decent - Mean scenario results in even higher LDES requirements (e.g., 1973/74) in which unusually mild conditions offset some of this leverage effect. Yet, in other years the year-specific heat demand leads to far higher LDES capacities, as VRE scarcities coincide with cold spells in a compound effect. While this effect only explains 25% on average, cold spells have a considerable impact in some of the most extreme weather years.
@@
LDES trajectories shown against daily deviations in heat demand from long-run mean (bright red - mid) and 24-hour rolling renewable availability (green - low). Gray area represents storage trajectory in No Heat. The darker red area shows the additional storage level in Decent - Mean, and the lighter red addition represents the incremental storage level from the specific weather year’s heat demand (Decent).
We illustrate both effects for four weather years with high additional LDES requirements in Figure 2. In the deterministic setting addressed here, the maximum required storage level over the course of a weather year determines the optimal LDES energy capacity. The trajectories therefore provide an insight into the critical periods that determine optimal LDES capacities. The gray areas correspond to the storage levels observed in the scenario without electrified heat (No Heat). The darker red area reflects additional storage levels required to meet the mean heat demand profile (Decent - Mean). The lighter red areas show the storage trajectories arising from the weather year’s specific heat demand profile (Decent). In the lower part of the figure, we plot the absolute daily deviations from long-run mean profiles of both heat demand (in dark-red) and renewable generation. The heat deviations serve as a proxy for cold spells. For renewable generation, we take the mean generation portfolio of the Decent scenario as a basis. The deviations are averaged in a 24-hour rolling window and are indicative of renewable energy scarcity, where the area under the zero line can be interpreted as an energy deficit.
@@
The weather year of 1958/59 (center right) results in the highest LDES capacity in the No Heat scenario due to a severe episode of energy scarcity between November and early January, leading to a discharge event of 79 TWh over the course of 28 days. Rather mild temperatures induce a LDES increase almost entirely through the leverage effect, where the discharge event increases in size to 223 TWh end extends to 106 days.
@@ 2.3 Long-duration thermal storage in district heating networks can mitigate LDES requirements
In contrast to a hydrogen cavern storage, a LDTS has relatively high standing losses from heat loss [sifnaios_heat_2023]. In the base specification, we assume an efficiency equivalent to cumulative losses of 39% within 90 days of storage in energetic terms [zeyen_mitigating_2021]. Hydrogen-based electricity storage, which includes electrolysis, cavern storage, and reconversion to electricity with gas turbines, has a relatively low roundtrip efficiency of around 30%, but is independent of the storage duration as standing losses are negligible. Even under these conservative estimates for LDTS efficiency, a LDTS system is more efficient in terms of electrical energy input per unit of thermal energy demand served for storage duration of up to ca. 180 days - longer than the seasonal heating cycle.
Our modeling results confirm such back-of-the-envelope calculations. The introduction of LDTS reduces the additional LDES requirements from electrified heat by 60 TWh (36%) on average (Figure 3, left panel). In the extreme weather year 1962/63 the introduction of co-optimized LDTS capacities in DH networks leads to a reduction of 155 TWh (37%). The Europe-wide average LDTS capacity over all years is 143 TWh, reaching almost 323 TWh in the maximum (Figure 3, center panel).
@@ 2.4 Central heat storage decentralizes overall long-duration storage infrastructure
Potentials for cavern storage capacities are unevenly distributed across Europe (Figure SI.10, left panel). Where there is potential in the form of rock salt formations in a country, the potential far exceeds any conceivably necessary storage requirement. Any of those countries could technically host all the cavern storage required for the entire continent. Yet, in the Decent scenario around 43% (72 out of 168 TWh) of the total cavern storage capacity is located in Germany on average, compare Figure SI.3. Notably, cavern storage is much more evenly distributed between countries in the No Heat scenario (Figure SI.2, left panel). Hydrogen gas turbines for reconversion are also concentrated in Germany in Decent, while electrolysis capacity is distributed more evenly, across countries on average.
Figure 4 shows the change in the geographical long-duration storage distribution across Europe when district heating networks with LDTS are introduced for the example of 1962/63. Heating demands and district heating potentials vary between countries such that a meaningful comparison of LDTS deployment requires a relative basis. We express LDTS as a share of total demand in district heating networks. LDTS uptake is particularly high in Central Eastern Europe where there is very little LDES capacity. The reduction in LDES capacity is disproportionally high in Germany. Consequently, the share of European LDES capacity in Germany falls to 35% (38 out of 108 TWh) in District & Decent. This masks two effects. For one, Germany has a relatively high district heating potential compared to other cavern storage locations, which means that it can substitute some of its on flexibility requirements locally. For two, German LDES provides flexibility in other countries, particularly in Central Eastern Europe and high LDTS adoption in these countries also reduces German LDES requirements.
@@ 2.5 Sensitivities
We test alternative specifications of LDTS efficiencies and investment costs. For each of the four weather years, we run 42 alternative settings, varying the LDTS standing losses on the x-axis and the investment costs on the y-axis (Figure 5). The base configuration used in all prior scenarios above is in the corner in the bottom-left. In the most favorable parameter configurations LDTS reduces LDES needs by another 19-35% (upper panel). In the case of 1941/42 this would correspond to an LDES requirement of 111 TWh when LDTS is free and efficiency is at 99%. As per the lower panel of Figure 5, such additional reductions are accompanied by large increases in optimal LDTS capacities. For 1941/42, the LDTS capacity almost triples between the base case and the best parameter combination. Based on the conclusion above that even in the base specification heat losses do not match the roundtrip losses of LDES for relevant storage durations, it is not surprising that additional substitution from efficiency gains beyond a threshold is limited. In the top line of each tile LDTS can be built for free. In these cases, the substitution potential with respect to LDES is limited by the country-level DH potentials.
@@
Sensitivity of LDTS parameters. The x-axis varies heat losses from the base case to almost no losses in energetic terms, expressed as energy remaining after 90 days in storage. The y-axis varies the assumed investment costs of LDTS by multiplying them with the labeled scalers.
@@
We also test the impact of different interconnection settings. First, we assume that there is no hydrogen network, see rows 3 and 7 of Table 2. Here, the increase in LDES energy capacities is more pronounced in the District & Decent scenario (20% - 44%) than in the Decent scenario (8% - 32%). In the District & Decent base case, hydrogen imports with long-term contracts are actually slightly higher than in the Decent base case, which is likely to replace some LDES capacity otherwise used for serving flat industrial hydrogen demand. In absence of hydrogen import capacities, this option no longer exists.
In the scenario with neither hydrogen nor electricity interconnection capacity (’Islands’, rows 4 and 8), LDES capacities increase only marginally compared to the ’No H2 network’ case and even decrease in the case of 1996/97 in the Decent scenario, indicating that the system value of LDES depends on interconnection. Without interconnection, building larger generation portfolios is cheaper than building expensive hydrogen tank storage. Only about 10 TWh of tank storage are being built across Europe in the ’Islands’ case of the Decent scenario.
Lastly, we test the effect of introducing around 97 GW of nuclear capacity, of which the majority is located in France (58 GW). In either scenario, it has a limited effect on LDES capacities (-14% and -18% on average). The reason is likely to be twofold. For one, nearly two thirds of the capacity being in France, limited interconnection capacities may prevent spatial balancing during extreme system episodes. By contrast, LDES leaves the option to transport both electricity and hydrogen. For two, as Section 2.3 shows, peak supply by hydrogen turbines in winter periods clearly exceeds the aggregate nuclear capacity in extreme years, and hence the introduced nuclear capacities can only supply a limited part of the positive residual load.
@@ 3.2 Limitations
The large geographic and temporal scope of our analysis comes at the expense of simplifications in the representation of several model components. Specifically, we only use one type of air-sourced heat pump for decentralized heating and do not include any efficiency improvements in the building stock. Similarly, we aggregate all district heating networks within a country and only consider one large-scale heat pump to supply these, abstracting from other potential heat sources. We further use a simplified representation of the thermodynamics of large-scale pit thermal storage, which neglects that some electricity might be required at the time of LDTS discharging to elevate the temperature level again, which could reduce the cost advantage of thermal storage against hydrogen storage. While some of these assumptions (e.g., only air-sourced heat pumps, no other heat sources) lead to a higher electricity demand and higher LDES needs than in reality expected, others might lead to an underestimation (simplified representation of the thermodynamics). Finally, the perfect foresight assumption for a given weather year leads to ideal but unrealistic storage operations and is likely to underestimate the relative value of LDES.
@@ 3.3 Conclusions & outlook
Long permitting processes and underdeveloped supply chains, as well as long-term uncertainty about hydrogen prices and quantities are likely to increase project risks [steinbach_future_2024]. Similarly, neither electrolysis nor hydrogen gas turbines have been deployed at scale so far [odenweller_green_2025]. Moreover, uncertainties around the ramp-up of heat electrification and hydrogen transmission infrastructure complicate the economics of highly centralized long-duration electricity storage (LDES). In this context, LDTS could provide a promising complementary technology, as it does not rely on particular geological conditions and can effectively deal with long-duration flexibility requirements in the heating sector. Yet, it is likely to be confined to heating networks and our results suggest that even if they were substantially expanded, a sizable residual long-duration flexibility need from the heating sector would remain.
Our results indicate that LDES in the form of cavern storage would be concentrated in a few countries, particularly in Germany, and thus other countries without salt cavern potentials to store hydrogen would be dependent on imports and transmission infrastructure to cover their load and heat demand. As this could raise regulatory and geopolitical questions, LDTS could for some countries increase their energy autonomy and decrease dependence on energy transport infrastructure.
@@
From an investor’s perspective, the increased variance of optimal LDES capacities between weather years when considering electrified heating raises additional questions around predictable cost recovery without other economic incentives. This clearly calls for de-risking such investments. Compared to hydrogen storage, it might be easier to manage such risks for thermal storage, as the latter is likely to be developed and operated by vertically integrated, cost-regulated regional monopolists who should be better able to manage quantity and price risks.
@@
Finally, the use of deterministic models omits a potentially important advantage of LDES over LDTS in settings without perfect foresight. LDES can serve final energy demand in many different sectors depending on the realized system state thanks to the usefulness and convertibility of hydrogen. In contrast, LDTS is less useful to the energy system as it can only serve heat demand. Future research, using stochastic modelling (such as in [schmidt_long-duration_2025]) could be a starting point to better understand the value of different types of storage, the interaction between them, and their operation in a non-deterministic setting. However, such approaches are likely to be feasible only for less extensive geographical and temporal scopes than modeled here.

### v2

@@ 1 Introduction
The direct or indirect electrification of sectors such as transport, industry, and heating is crucial for the complete decarbonization of economies around the world [international_energy_agency_net_2024]. Yet, most studies at the nexus of LDES and weather variability focus only on the power sector and neglect the impact of weather variations on the demand side of a sector-coupled renewable system [gotske_designing_2024]. Accounting for 16% of total greenhouse gas emissions in the European Union in 2024 [rozsai_m_jrc-idees-2021_2024], the heating sector takes center stage in the continent’s decarbonization efforts [commission_renewable_2023]. Previous studies have shown that serving a substantial part of residential and commercial heat demand with heat pumps can have challenging consequences for the power system [bloess_power--heat_2018, zeyen_mitigating_2021, roth_power_2023, roth_power_2024]. Furthermore, there is a substantial degree of interannual variation in temperatures, and correspondingly in heat demand and temperature-dependent heat pump efficiencies. There is evidence that cold spells can coincide with times of renewable supply shortages [van_der_wiel_influence_2019, mockert_meteorological_2023] exacerbating system stress.
@@ 2 Results
In the following, we focus primarily on the link between the electrification of residential and commercial heat and LDES needs. The Supplemental Information Section SI.4 provides several additional results. We summarize all considered scenarios in Table 1 and provide additional details in Section 4.3.
@@ 2.1 Electrified heat demand leads to significantly higher long-duration electricity storage requirements
In Figure 1, the weather years are sorted by LDES requirement in the Decent scenario. Notably, the No Heat scenario does not reflect the same order. Weather years that have high LDES requirements in the scenario without electrified heat do not necessarily rank high when considering electrified heat (cf. 1989/90) and vice versa (cf. 1941/42). Additionally, the distribution of optimal LDES capacities widens substantially. In the No Heat scenario, the standard deviation of total LDES energy capacities across all weather years is 10 TWh. In the scenario Decent, the standard deviation increases to 66 TWh.
@@
Optimal long-duration electricity storage capacities by weather year. Gray dots represent optimal capacities in the No Heat scenario. Dark-red dots represent optimal capacities in the Decent scenario. Light-red crosses represent optimal capacities in the Decent - Mean scenario, in which weather-year specific heat demand is replaced by long-run mean heat demand.
@@ 2.2 Heat demand drives storage needs through ‘leverage’ and ‘compound’ effects
In a third set of scenario runs (Decent - Mean), we replace the time series of heat demand and heat pump coefficients of performance (COP) in a given weather year by their long-run hourly averages across all weather years in the sample. These curves (see Figure SI.11) isolate the inherent seasonality of electricity demand for heating and abstract from cold spells or unusually warm periods. The results of these model runs (light red crosses in Figure 1) highlight the drivers of increasing LDES requirements. Heat demand is seasonal, and electrified heating naturally adds most load in winter months. As extreme system states of VRE scarcity mostly occur during winter months [kittel_quantifying_2024, kittel_coping_2025, gotske_designing_2024], the additional load amplifies the energy deficit, triggering additional LDES requirements by functioning as a lever. Across the sample of weather years, this leverage effect accounts for 98 TWh or 75% of the total LDES increase on average.
For some years, the Decent - Mean scenario results in even higher LDES requirements (e.g., 1973/74) for which unusually mild conditions offset some of this leverage effect. Yet, in other years, the year-specific heat demand leads to far higher LDES capacities, as VRE scarcities coincide with cold spells in a compound effect. While this effect only explains 25% on average, cold spells have a considerable impact in some of the most extreme weather years.
@@
LDES trajectories shown against daily deviations in heat demand from long-run mean (bright red - mid) and 24-hour rolling renewable availability (green - low). The gray area represents the storage trajectory in No Heat. The darker red area shows the additional storage level in Decent - Mean, and the lighter red addition represents the incremental storage level from the specific weather year’s heat demand (Decent).
We illustrate both effects for four weather years with high additional LDES requirements in Figure 2. In the deterministic setting addressed here, the maximum required storage level in the course of a weather year determines the optimal LDES energy capacity. The trajectories, therefore, provide an insight into the critical periods that determine optimal LDES capacities. The gray areas correspond to the storage levels observed in the scenario without electrified heat (No Heat). The darker red area reflects additional storage levels required to meet the mean heat demand profile (Decent - Mean). The lighter red areas show the storage trajectories arising from the weather year’s specific heat demand profile (Decent). In the lower part of the figure, we plot the absolute daily deviations from long-run mean profiles of both heat demand (in dark red) and renewable generation. The heat deviations serve as a proxy for cold spells. For renewable generation, we take the mean generation portfolio of the Decent scenario as a basis. The deviations are averaged in a 24-hour rolling window and are indicative of renewable energy scarcity, where the area under the zero line can be interpreted as an energy deficit.
@@
The weather year of 1958/59 (center right) results in the highest LDES capacity in the No Heat scenario due to a severe episode of energy scarcity between November and early January, leading to a discharge event of 79 TWh over the course of 28 days. Rather mild temperatures induce a LDES increase almost entirely through the leverage effect, where the discharge event increases in size to 223 TWh and extends to 106 days.
@@ 2.3 Long-duration thermal storage in district heating networks can mitigate LDES requirements
In contrast to a hydrogen cavern storage, an LDTS has relatively high standing losses from heat loss [sifnaios_heat_2023]. In the base specification, we assume an efficiency equivalent to cumulative losses of 39% within 90 days of storage in energetic terms [zeyen_mitigating_2021]. Hydrogen-based electricity storage, which includes electrolysis, cavern storage, and reconversion to electricity with gas turbines, has a relatively low roundtrip efficiency of around 30%, but is independent of the storage duration as standing losses are negligible. Even under these conservative estimates for LDTS efficiency, a LDTS system is more efficient in terms of electrical energy input per unit of thermal energy demand served for storage duration of up to ca. 180 days - longer than the seasonal heating cycle.
Our modeling results confirm such back-of-the-envelope calculations. The introduction of LDTS reduces the additional LDES requirements from electrified heat by 60 TWh (36%) on average (Figure 3, left panel). In the extreme weather year 1962/63, the introduction of co-optimized LDTS capacities in DH networks leads to a reduction of 155 TWh (37%). The Europe-wide average LDTS capacity over all years is 143 TWh, reaching almost 323 TWh in the maximum (Figure 3, center panel).
@@ 2.4 Central heat storage decentralizes overall long-duration storage infrastructure
Potentials for cavern storage capacities are unevenly distributed across Europe (Figure SI.10, left panel). Where there is potential in the form of rock salt formations in a country, the potential far exceeds any conceivably necessary storage requirement. Any of those countries could technically host all the cavern storage required for the entire continent. Yet, in the Decent scenario, around 43% (72 out of 168 TWh) of the total cavern storage capacity is located in Germany on average, compare Figure SI.3. Notably, cavern storage is much more evenly distributed between countries in the No Heat scenario (Figure SI.2, left panel). Hydrogen gas turbines for reconversion are also concentrated in Germany in Decent, while electrolysis capacity is distributed more evenly, across countries on average.
Figure 4 shows the change in the geographical long-duration storage distribution across Europe when district heating networks with LDTS are introduced for the example of 1962/63. Heating demands and district heating potentials vary between countries, such that a meaningful comparison of LDTS deployment requires a relative basis. We express LDTS as a share of total demand in district heating networks. LDTS uptake is particularly high in Central Eastern Europe where there is very little LDES capacity. The reduction in LDES capacity is disproportionately high in Germany. Consequently, the share of European LDES capacity in Germany falls to 35% (38 out of 108 TWh) in District & Decent. This masks two effects. For one, Germany has a relatively high district heating potential compared to other cavern storage locations, which means that it can substitute some of its on flexibility requirements locally. For another, German LDES provides flexibility in other countries, particularly in Central Eastern Europe and high LDTS adoption in these countries also reduces German LDES requirements.
@@ 2.5 Sensitivities
We test alternative specifications of LDTS efficiencies and investment costs. For each of the four weather years, we run 42 alternative settings, varying the LDTS standing losses on the x-axis and the investment costs on the y-axis (Figure 5). The base configuration used in all prior scenarios above is in the bottom-left corner. In the most favorable parameter configurations LDTS reduces LDES needs by another 19-35% (upper panel). In the case of 1941/42, this would correspond to an LDES requirement of 111 TWh when LDTS is free and efficiency is at 99%. As per the lower panel of Figure 5, such additional reductions are accompanied by large increases in optimal LDTS capacities. For 1941/42, the LDTS capacity almost triples between the base case and the best parameter combination. Based on the conclusion above that even in the base specification, heat losses do not match the round-trip losses of LDES for relevant storage durations, it is not surprising that additional substitution from efficiency gains beyond a threshold is limited. In the top line of each tile, LDTS can be built for free. In these cases, the substitution potential with respect to LDES is limited by the country-level DH potentials.
@@
Sensitivity of LDTS parameters. The x-axis varies heat losses from the base case to almost no losses in energetic terms, expressed as energy remaining after 90 days in storage. The y-axis varies the assumed investment costs of LDTS by multiplying them by the labeled scalers.
@@
We also test the impact of different interconnection settings. First, we assume that there is no hydrogen network, see rows 3 and 7 of Table 2. Here, the increase in LDES energy capacities is more pronounced in the District & Decent scenario (20% - 44%) than in the Decent scenario (8% - 32%). In the District & Decent base case, hydrogen imports with long-term contracts are actually slightly higher than in the Decent base case, which is likely to replace some LDES capacity otherwise used for serving flat industrial hydrogen demand. In the absence of hydrogen import capacities, this option no longer exists.
In the scenario with neither hydrogen nor electricity interconnection capacity (’Islands’, rows 4 and 8), LDES capacities increase only marginally compared to the ’No H2 network’ case and even decrease in the case of 1996/97 in the Decent scenario, indicating that the system value of LDES depends on interconnection. Without interconnection, building larger generation portfolios is cheaper than building expensive hydrogen tank storage. Only about 10 TWh of tank storage is being built across Europe in the ’Islands’ case of the Decent scenario.
Lastly, we test the effect of introducing around 97 GW of nuclear capacity, of which the majority is located in France (58 GW). In either scenario, it has a limited effect on LDES capacities (-14% and -18% on average). The reason is likely to be twofold. For one, nearly two-thirds of the capacity is in France, and limited interconnection capacities may prevent spatial balancing during extreme system episodes. By contrast, LDES leaves the option to transport both electricity and hydrogen. For two, as Section 2.3 shows, peak supply by hydrogen turbines in winter periods clearly exceeds the aggregate nuclear capacity in extreme years, and hence the introduced nuclear capacities can only supply a limited part of the positive residual load.
@@ 3.2 Limitations
The large geographic and temporal scope of our analysis comes at the expense of simplifications in the representation of several model components. Specifically, we only use one type of air-sourced heat pump for decentralized heating and do not include any efficiency improvements in the building stock. Similarly, we aggregate all district heating networks within a country and only consider one large-scale heat pump to supply these, abstracting from other potential heat sources. We further use a simplified representation of the thermodynamics of large-scale pit thermal storage, which neglects that some electricity might be required at the time of LDTS discharging to elevate the temperature level again, which could reduce the cost advantage of thermal storage against hydrogen storage. While some of these assumptions (e.g., only air-sourced heat pumps, no other heat sources) lead to a higher electricity demand and higher LDES needs than expected in reality, others might lead to an underestimation (simplified representation of the thermodynamics). Finally, the perfect foresight assumption for a given weather year leads to ideal but unrealistic storage operations and is likely to underestimate the relative value of LDES.
@@ 3.3 Conclusions & outlook
Long permitting processes and underdeveloped supply chains, as well as long-term uncertainty about hydrogen prices and quantities, are likely to increase project risks [steinbach_future_2024]. Similarly, neither electrolysis nor hydrogen gas turbines have been deployed at scale so far [odenweller_green_2025]. Moreover, uncertainties around the ramp-up of heat electrification and hydrogen transmission infrastructure complicate the economics of highly centralized long-duration electricity storage (LDES). In this context, LDTS could provide a promising complementary technology, as it does not rely on particular geological conditions and can effectively deal with long-duration flexibility requirements in the heating sector. Yet, it is likely to be confined to heating networks, and our results suggest that even if they were substantially expanded, a sizable residual long-duration flexibility need from the heating sector would remain.
Our results indicate that LDES in the form of cavern storage would be concentrated in a few countries, particularly in Germany, and thus other countries without salt cavern potentials to store hydrogen would be dependent on imports and transmission infrastructure to cover their load and heat demand. As this could raise regulatory and geopolitical questions, LDTS could, for some countries, increase their energy autonomy and decrease dependence on energy transport infrastructure.
@@
From an investor’s perspective, the increased variance of optimal LDES capacities between weather years, when considering electrified heating, raises additional questions around predictable cost recovery without other economic incentives. This clearly calls for de-risking such investments. Compared to hydrogen storage, it might be easier to manage such risks for thermal storage, as the latter is likely to be developed and operated by vertically integrated, cost-regulated regional monopolists who should be better able to manage quantity and price risks.
@@
Finally, the use of deterministic models omits a potentially important advantage of LDES over LDTS in settings without perfect foresight. LDES can serve final energy demand in many different sectors, depending on the realized system state, thanks to the usefulness and convertibility of hydrogen. In contrast, LDTS is less useful to the energy system as it can only serve heat demand. Future research, using stochastic modeling (such as in Schmidt [schmidt_long-duration_2025]) could be a starting point to better understand the value of different types of storage, the interaction between them, and their operation in a non-deterministic setting. However, such approaches are likely to be feasible only for less extensive geographical and temporal scopes than modeled here.